OpenStudy (anonymous):

1.Let g(x) =root(x^2 + 16 ) a.Find the linearization of g(x) at x = 3. b.Use your linearization to approximate g(2.9).

OpenStudy (j-cross18):

Hay are you still there???

OpenStudy (anonymous):

yes

OpenStudy (j-cross18):

ok I do not know what to do sorry lol but if you want you could give me a medal for trying lol PLz ^.^ Thanks

OpenStudy (j-cross18):

Can you please @butterflyprincess ????

OpenStudy (anonymous):

use the formula L(x) = g(a) + g'(a)(x-a) part a: when they say at x=3 they want the tangent line at the point (3, 5) so its a = 3 when finding the linearization \[g(3) = \sqrt{3^2 + 16} = 5\] g'(x) = \[\frac{ 1 }{ 2 } (x^2 + 16)^\frac{ -1 }{ 2 } (2x) = \frac{ x }{ \sqrt{x^2 + 16} }\] so g'(3) = 3/5 so now you have L(x) = 5 + 3/5(x-3) L(x) = 5 + 3/5x - 9/5 \[L(x) = \frac{ 16 }{ 5 } + \frac{ 3 }{ 5 }x \] now for part b: to estimate g(2.9) so substitute it in the linearization equation L(2.9) = 16/5 + 3/5(2.9) = 4.94 you can check this by doing g(2.9) = 4.94064773